A one-meter rod of 2 cm diameter is drawn until its resistance is 100 times the initial resistance. Its length afterward is?


1
Expert's answer
2021-10-06T00:31:27-0400

For its initial state, write how resistance depends on other characteristics:


"R=\frac{\rho L}{A}=\frac{4\rho L}{\pi D^2}."

For the drawn state:


"R_d=100R=\frac{4\rho L_d}{\pi d^2}."

Since we know the initial volume of the material, find the diameter of the thin wire:


"V=L\frac{\pi D^2}{4}=L_d\frac{\pi d^2}{4},\\\space\\
\pi d^2=\frac{L\pi D^2}{L_d}."

Substitute this into the equation for Rd:

"100R=\frac{4\rho L_d^2}{L\pi D^2},\\\space\\
100·\frac{4\rho L}{\pi D^2}=\frac{4\rho L_d^2}{L\pi D^2},\\\space\\
100·L=\frac{L_d^2}{L},\\\space\\
L_d=10L=10\text{ m}."